OpenStudy (anonymous):

verify folllowing identities sin3x-sinx=2cos2xsinx

OpenStudy (anonymous):

sin(3x)-sin(x)=sin(2x+x)-sin(x)= sin2xcosx+sinxcos2x-sinx=2sinxcos^2x+sinxcos2x-six= 2sinxcos^2x-sinx+sinxcos2x=sinx(2cos^2x-1)+sinxcos2x=sinxcos2x+sinxcos2x= 2sinxcos2x